OpenStudy (anonymous):

If f(x) = 4x + 8 and g(x) = sqrt x + 4, what is (f o g)(12)? Help please.. So confused.

OpenStudy (tkhunny):

\(f(g(x)) = f\left(\sqrt{x+4}\right) = 4\left(\sqrt{x+4}\right) + 8\) What else would it be?

OpenStudy (anonymous):

I don't know??

OpenStudy (tkhunny):

Okay, then if you don't know what else it would be, why not go with what you were just shown? It's just substitution. That's all that is going on.

OpenStudy (anonymous):

Okay, so where do I go from there?

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@Loser66 help?

OpenStudy (loser66):

tkhunny gave you the whole answer, where do you want to go ? DAT SIT

OpenStudy (tkhunny):

f(x) = 3x - 5 f(2) = 3(2) - 5 f(Dog) = 3(Dog) - 5 f(g(x)) = 3(g(x)) - 5 Substitution.